# Adaptive Reduced Tempering For Bayesian inverse problems and rare event simulation

### v1

@@ Abstract
This work proposes an adaptive sequential Monte Carlo sampling algorithm for solving inverse Bayesian problems in a context where a (costly) likelihood evaluation can be approximated by a surrogate, constructed from previous evaluations of the true likelihood. A rough error estimation of the obtained surrogates is required. The method is based on an adaptive sequential Monte-Carlo (SMC) simulation that jointly adapts the likelihood approximations and a standard tempering scheme of the target posterior distribution. This algorithm is well-suited to cases where the posterior is concentrated in a rare and unknown region of the prior. It is also suitable for solving low-temperature and rare-event simulation problems. The main contribution is to propose an entropy criteria that associates to the accuracy of the current surrogate a maximum inverse temperature for the likelihood approximation. The latter is used to sample a so-called snapshot, perform an exact likelihood evaluation, and update the surrogate and its error quantification. Some consistency results are presented in an idealized framework of the proposed algorithm. Our numerical experiments use in particular a reduced basis approach to construct approximate parametric solutions of a partially observed solution of an elliptic Partial Differential Equation. They demonstrate the convergence of the algorithm and show a significant cost reduction (close to a factor $10$) for comparable accuracy.
@@ 1.1 Context and problem
We consider the classical problem of designing an algorithmic Monte Carlo procedure that samples according to a target probability distribution $\eta^{\ast}_{\beta_{\infty}}$ (for one specific $\beta_{\infty}\in\mathbb{R}$) defined explicitly up to a normalization as an element of the family
@@
In the above, the quantity $Z^{\ast}_{\beta}\displaystyle\mathop{=}^{{\rm def}}\pi(e^{\beta S^{\ast}})$ denotes the associated unknown normalization which typically must also be computed by the Monte Carlo procedure, $\pi$ denotes a given, easily simulable, reference probability distribution over a state space $\mathcal{X}$, and $S^{\ast}:\mathcal{X}\to\mathbb{R},$ denotes a given yet expensive to compute real-valued score function.
The following two contexts will be used as the main motivations. In both cases, the score function is of the form
@@
  • In the Bayesian formulation of inverse problems, $\beta_{\infty}=1$ and $\eta^{\ast}_{\beta_{\infty}}$ is the posterior distribution of some uncertainty parameters $x\in\mathcal{X}$ of a partially observed physical system $\Psi^{\ast}(x)$. $\pi$ models the prior distribution of the uncertainty parameters and the weight $\propto e^{S^{\ast}}$ is the likelihood of those parameters given some partial observations. As usual, one wants to Monte Carlo sample and then approximate expectations according to this posterior.
  • In rare events problems, one wants to estimate the small probability $p^{\ast}\displaystyle\mathop{=}^{{\rm def}}\pi\left(\left\{S^{\ast}\geq 1\right\}\right)$ corresponding to a large prescribed level $\ell\in\mathbb{R}$, with a score defined as $S^{\ast}(x)=1-\max(\ell-\Psi^{\ast}(x),0)/\ell$. As $\lim_{\beta_{\infty}\to\infty}\pi(e^{\beta_{\infty}S^{\ast}})=p^{\ast}$, the estimation of the rare event probability turns out to be the estimation of the normalization constant $Z^{\ast}_{\beta_{\infty}}$ as $\beta_{\infty}\to\infty$, i.e., as the temperature $1/\beta_{\infty}$ tends to zero.
In this work, we will be interested in scenarios in which the practitioner has to face two types of difficulty. The first difficulty is classical and arises when $\beta$ increases (or when the measurement noise becomes small in a Bayesian context), yielding a low-temperature problem, similar to global optimization. It confronts the practitioner with a computational challenge: the density defining $\eta^{\ast}_{\beta}$ concentrates the majority of the probability mass in specific but possibly unknown areas of $\mathcal{X}$, which are roughly described by the maxima of $S^{\ast}$. A direct sampling with a Monte Carlo simulation of $\pi$ is often time-consuming or even infeasible, as it can require a very large sample size. For instance to obtain a reasonable estimation variance, the sample size should be of order of $1/p^{\ast}$ in the rare event simulation context, which is obviously prohibitive. A very popular general strategy to simulate a sample according to $\eta^{\ast}_{\beta}$ is to resort to a Sequential Monte Carlo (SMC) strategy (see [16]) that starts with a Monte Carlo sample of particles distributed according to $\pi$ and then sequentially samples $\eta^{\ast}_{\beta}$ for increasing values of $\beta$ by resorting to a combination of Importance Sampling (IS), re-sampling (selection) of weighted samples and mutations of particles based on suitable Markov Chain Monte Carlo transitions. More specifically, we will be interested in the adaptive variant of tempering [3].
@@
The first feature is an a posteriori pointwise error quantification associated with $S$. This error quantification comes in the form of function $E(x)$ satisfying an approximate pointwise estimate of the form
@@
We underline that the pointwise estimate (1.2) need not to be a very precise estimate on the error, but only a rough estimate revealing the trend of the error.
@@
[…]
@@
The associated error estimate in (1.2) is updated simultaneously, and for reasons that will be explained in Section 2.2, we additionally require that the error $E^{(k)}$ vanishes on snapshot points $X^{(k)}$ already evaluated at iteration $k$ :
$E^{(X_{1},\ldots,X_{k})}(X^{(j)})=0,\quad\forall j\leq k.$ (1.3)
In order to improve the readability, and throughout the paper, we may use the superscript $(k)$ instead of $(X_{1},\ldots,X_{k})$, or even drop it for any quantity that depends on the sample of snapshots $(X_{1},\ldots,X_{k})$. Fo instance the reduced score constructed out of $k$ snapshots may be denoted:
@@ 1.2 Previous works
Advanced methods for Bayesian sampling of (1.1) are gradient-based Markov chain Monte Carlo (MCMC) methods such as Hamiltonian Monte Carlo [6] and preconditioning (of gradients) techniques [14]. Estimation with such advanced MCMC techniques can still exhibit slow convergence rates, as the simulation may need a large time to explore the target distribution, especially in high-dimension [22]. Sequential Monte Carlo (SMC) methods or other methods based on interacting clones/particles are an interesting alternative for Bayesian sampling, and have shown to be stable in certain high-dimensional context [2], in particular adaptive SMC simulation [3]. These methods have been employed to solve Bayesian inverse problems [5, 23]. Other sampling methods such as parallel tempering have been proposed to accelerate convergence [24]. Nevertheless, sampling with any of these methods become difficult in the present context, where only a too small budget of evaluations of the score function $S^{\ast}$ is affordable.
@@
In recent years, numerous MCMC and SMC methods based on reduced score $S$ approximating $S^{\ast}$ have been studied to reduce the simulation cost. Among MCMC approaches for inverse Bayesian problems, as expected, convergence results have been obtained for a sequence of reduced scores converging to $S^{\ast}$ [13]. Refinements of this approach involve integrating an iterative update of the $S^{\ast}-S$ approximation error into a reduced MCMC sampling scheme using $S$, for a given a predetermined estimate of the error (based on the evaluation of the error for a set of samples drawn from the prior) [7]. It should be noted, however, that a priori sampling to build the reduced model, characterize the approximation error or do both is probably not a relevant approach in our context. Indeed, in Bayesian inverse problems with informative measurements or in low-temperature problems, the target distribution usually concentrates its mass in a small and specific but unknown region of the prior support. More adapted to this context is the strategy to adapt the reduced model during the simulation, using samples whose distribution gradually matches the final target distribution.
@@
Adaptive construction of the reduced score embedded in a tempering scheme was suggested in [26]. As mentioned above, such an adaptive construction of the approximation is relevant to our context, but this work remains quite specific as it relies on parametric Gaussian proposals designed by a cross-entropy method. Moreover, the adaptive construction of reduced score does not use possible error estimates.
@@
Regarding SMC approaches, multilevel SMC has been proposed to optimize the product cost $\times$ variance of the estimates; in the latter ’multilevel’ refers to a sequence of approximations of the score function at different discretization levels [4]. However, multilevel SMC requires precise characterization of the model-dependent variance bounds of the estimates, which is a tedious task. In addition, although it relies on a given hierarchy of reduced models or approximations, these are not updated online using snapshots.
@@
An adaptive SMC approach using a predetermined hierarchy of reduced scores has also been proposed in [25]. This work tackles the tricky problem of selecting the right reduced model at the current temperature during simulation. Indeed, choosing a proper criterion to decide along the simulation when to refine the reduced score is quite an open question in general [20], although various heuristics have been proposed in the context or rare event simulation relying on some error quantification [1, 21, 11]. In their work, the authors of [25] propose a heuristic consisting on minimizing the variation of the particules weights testing the case of a reduced score refined or not. In this work, we go beyond these ideas and generally answer the question of when to update the reduced score and with which snapshots.
@@
The authors of [25] have also proposed an analogous method [30] in the context of rare event simulation, in which a predetermined hierarchy of reduced model is given. The authors of the present work have also proposed a method similar to the one studied in the present paper in a rare event context: the adaptive reduced multilevel splitting (ARMS) algorithm [11] (Nota Bene: ’multilevel’ here has historically a completely different meaning as the one in ’multilevel SMC’ mentioned before). ARMS is based on a sequential simulation scheme very similar to SMC, which adapts a reduced score approximation while getting closer to the target rare event distribution. This method is nevertheless specific to the rare event simulation context in the sense it is dedicated to sample targets being the reference distribution conditionally to be in some rare event set. The method proposed in this paper (adaptive reduced tempering, ART) can be seen as a generalization of the ARMS algorithm to the case of general target distributions of the form of (1.1), which are relevant for Bayesian problems. However, as explained e.g. in Section 1.1, the ART method is also suited to the context of rare event simulation, by considering the latter as a zero temperature limit of a Gibbs family.
As compared to the pieces of work mentioned above, the main novelty of the proposed ART algorithm is to propose an adaptive rule that triggers, within a SMC sampling method with tempered surrogate targets, when to sample a new snapshot of the true model (serving to update of the reduced model). Details are given in the next section.
@@ 1.3 Contributions
This work proposes a methodology for sampling the flow of targets (1.1) combining an adaptive Sequential Monte Carlo approach and the availability of reduced modeling with error quantification. To do so, we consider the following exponential family of importance distributions
@@
where the normalization $Z_{\beta}\displaystyle\mathop{=}^{{\rm def}}\pi(\mathrm{e}^{\beta S})$ represents an approximation of $Z^{\ast}_{\beta}$. An important idea is to interpret the latter as proposal distributions (also called importance distributions), in the sense of importance sampling, of the target family (1.1) . In the context of rare event simulation, using proposal distributions in this exponential family rather than e.g. in the form of indicator of events $\propto\,{\rm l}{\rm l}_{S\geq\ell}\mathrm{d}\pi$ with $\ell\in\mathbb{R}$ enables to get rid of the possible nesting conditions between successive proposals in the algorithm (see [11]).
@@
where the reduced score $S^{(k)}$ is much cheaper to computes, and only the evaluation of the importance weight is expensive. Very importantly, as the iterations progress, the proposal will adapt by adjusting the inverse temperature $\beta^{(k)}$ and refining the approximation of the reduced score $S^{(k)}$ using the history of the previous computations of the true model. In this framework, three issues emerge.
A first important one is that one has to address the definition of a criterion for setting the lowest inverse temperature $\beta^{(k)}$ suitable for the accuracy of the current reduced score approximation $S^{(k)}$. Justified by theoretical arguments, we propose to constrain the relative entropy111 or Kullback-Leibler divergence between probabilities defined by $\mathrm{Ent}(\eta\mid\mu)\displaystyle\mathop{=}^{{\rm def}}\int\log\frac{\mathrm{d}\eta}{\mathrm{d}\mu}\mathrm{d}\eta$ when $\eta$ is dominated by $\mu$ and extended to +$\infty$+ ∞ otherwise. between a surrogate target denoted $\check{\eta}_{\beta}$ and the proposal as follows (for a given threshold $c_{1}$):
@@
The surrogate target is constructed using the approximate error bound (1.2) of the reduced score approximation as follows:
@@
The choice is motivated by the fact that $\check{\eta}_{\beta}$ is roughly speaking a version of $\eta_{\beta}$ where typical states with large expected errors (hence ’far away’ from the already sampled snapshots) are penalized.
A second issue that we will address in this work is the tuning of the expensive snapshot sampling and its associated reduced modeling update. At iteration $k+1$ the snapshot $X_{k+1}$ will be sampled from the proposal $\mu^{(k)}_{\beta^{(k)}}$ which ensures thanks to (1.5) a sampling in areas consistent with the true target distribution $\eta^{\ast}_{\beta^{(k)}}$ at inverse temperature $\beta^{(k)}$. In a first phase of the algorithm, we sample snapshots with an additional strong penalty weight favoring candidate states with a larger error (hence requiring priority correction). In a final phase of the algorithm, we remove that penalty to obtain snapshot-based optimal importance sampling estimation.
A last issue that has been addressed in this work is the simulation of a new proposal when at the end of iteration $k$ the reduced score is updated $S^{(k)}\to S^{(k+1)}$. We will propose to bridge proposals by searching among past simulations for an iteration index $k_{bridge}$ with associated approximate score $S^{(k_{bridge})}$, and by finding an inverse temperature $\beta_{bridge}$ such that the relative entropy between the new proposal $\mu_{\beta_{\mathrm{bridge}}}^{(k+1)}$ and the past proposal ${\mu_{\beta^{(k_{\mathrm{bridge}})}}^{(k_{\mathrm{bridge}})}}$ is kept below a certain standard threshold:
@@
This work will demonstrate the following facts about the proposed algorithmic methodology. First, at a mathematical level, the proposed importance sampling estimator is shown to be consistent in an idealized setting. Second, we exhibit in our numerical experiments its practical convergence. Finally and most importantly, by evaluating the computational complexity in relation to the estimation error, we show that the use of reduced modeling (as compared to similar methods with only true model evaluation) significantly reduces the simulation cost while achieving a comparable accuracy.
The paper is organized as follows. An idealized version of the proposed algorithm, in which we assume the exact evaluation of expectations and the exact sampling of distributions, is presented in Section 2 together with the underlying methodological concepts. We then provide in Section 3 a practical implementation of this idealized algorithm using SMC simulations. In Section 4, we numerically evaluate the proposed algorithm in the case of rare event simulation and inverse Bayesian problems. A final section deals with conclusions.

### v2

@@ Abstract
This work proposes an adaptive sequential Monte Carlo sampling algorithm to solve Bayesian inverse problems in scenarios where likelihood evaluations are costly but can be approximated using a surrogate model built from previous evaluations of the true likelihood. A rough estimate of the surrogate error is required. The method relies on an adaptive SMC framework that simultaneously adjusts both the likelihood approximations and a standard tempering scheme of the target posterior distribution. This algorithm is well-suited for cases where the posterior is concentrated in a rare and unknown region of the prior. It is also suitable for solving low-temperature and rare event simulation problems. The main contribution is to propose an entropy criterion that relates the accuracy of the current surrogate to a maximum inverse temperature for the likelihood approximation. The latter is instrumental to sample a so-called snapshot, on which is performed an exact likelihood evaluation, used to update the surrogate and its error quantification. Some consistency results are presented in an idealized framework for the proposed algorithm. Our numerical experiments use in particular a reduced basis approach to construct approximate parametric solutions to a partially observed solution of an elliptic partial differential equation. They demonstrate the convergence of the algorithm and show a significant cost reduction (close to a factor of $10$) for comparable accuracy.
@@ 1.1 Context and problem
We consider the classical problem of designing an algorithmic Monte Carlo procedure that samples from a target probability distribution $\eta^{\ast}_{\beta_{\infty}}$ (for one specific $\beta_{\infty}\in\mathbb{R}$) defined explicitly up to a normalization as an element of the family
@@
In the above, the quantity $Z^{\ast}_{\beta}\displaystyle\mathop{=}^{{\rm def}}\pi(e^{\beta S^{\ast}})$ denotes the associated unknown normalization which typically must also be computed by the Monte Carlo procedure. $\pi$ denotes a given, easily simulable, reference probability distribution over a state space $\mathcal{X}$, and $S^{\ast}:\mathcal{X}\to\mathbb{R}$ is a given but expensive-to-compute real-valued score function.
Two motivating contexts are considered, where the score function takes the form:
@@
  • In the Bayesian formulation of inverse problems, $\beta_{\infty}=1$ and $\eta^{\ast}_{\beta_{\infty}}$ is the posterior distribution of some uncertainty parameters $x\in\mathcal{X}$ in a partially observed physical system $\Psi^{\ast}(x)$. $\pi$ models the prior distribution of the uncertainty parameters and the weight $\propto e^{S^{\ast}}$ is the likelihood of those parameters given some partial observations. As usual, one wants to Monte Carlo sample and then approximate expectations with respect to this posterior.
  • In rare events problems, one wants to estimate the small probability $p^{\ast}\displaystyle\mathop{=}^{{\rm def}}\pi(\{S^{\ast}\geq 1\})$ corresponding to a large prescribed level $\ell\in\mathbb{R}$, with a score defined as $S^{\ast}(x)=1-\max(\ell-\Psi^{\ast}(x),0)/\ell$. As $\lim_{\beta_{\infty}\to\infty}\pi(e^{\beta_{\infty}S^{\ast}})=p^{\ast}$, the estimation of the rare event probability turns out to be the estimation of the normalization constant $Z^{\ast}_{\beta_{\infty}}$ as $\beta_{\infty}\to\infty$, i.e., as the temperature $1/\beta_{\infty}$ tends to zero.
In this work, we will be interested in scenarios in which the practitioner has to face two types of difficulty. The first difficulty is classical and arises when $\beta$ increases (or when the measurement noise becomes small in a Bayesian context), yielding a low-temperature problem, similar to global optimization. It confronts the practitioner with a computational challenge: the density defining $\eta^{\ast}_{\beta}$ concentrates the majority of the probability mass in specific but possibly unknown areas of $\mathcal{X}$, which are roughly described by the maxima of $S^{\ast}$. Direct sampling with a Monte Carlo simulation of $\pi$ is often time-consuming or even infeasible, as it can require a very large sample size. For instance to obtain a reasonable estimation variance, the sample size should be of order of $1/p^{\ast}$ in the rare event simulation context, which is obviously prohibitive. A very popular general strategy to simulate a sample according to $\eta^{\ast}_{\beta}$ is to resort to a Sequential Monte Carlo (SMC) strategy (see [19]) that starts with a Monte Carlo sample of particles distributed according to $\pi$ and then sequentially samples $\eta^{\ast}_{\beta}$ for increasing values of $\beta$ by resorting to a combination of Importance Sampling (IS), re-sampling (selection) of weighted samples and mutations of particles based on suitable Markov Chain Monte Carlo transitions. More specifically, we will be interested in the adaptive variant of tempering [5].
@@
The first feature is an a posteriori point-wise error quantification associated with $S$. This error quantification comes in the form of function $E(x)$ satisfying an approximate point-wise estimate of the form
@@
We underline that the point-wise estimate (1.2) need not to be a very precise estimate on the error, but only a rough estimate revealing the trend of the error.
@@
$\left\{(X_{1},\Psi^{\ast}(X_{1})),\ldots,(X_{k},\Psi^{\ast}(X_{k})\right\}\xmapsto{\textrm{reduced model}}\Psi=\Psi^{(X_{1},\ldots,X_{k})},$
@@
The associated error estimate in (1.2) is updated simultaneously, and for reasons that will be explained in Section 2.2, we additionally require that the error $E_{1},\ldots,E_{k}$ vanishes on snapshot points $X_{1},\ldots,X_{k}$ already evaluated at iteration $k$ :
$E^{(X_{1},\ldots,X_{k})}(X_{j})=0,\quad\forall j\leq k.$ (1.3)
In order to improve the readability, and throughout the paper, we may use the superscript $(k)$ instead of $(X_{1},\ldots,X_{k})$, or even drop it for any quantity that depends on the sample of snapshots $(X_{1},\ldots,X_{k})$. For instance the reduced score constructed out of $k$ snapshots may be denoted:
@@ 1.2 Previous works
Advanced methods for Bayesian sampling of (1.1) are gradient-based Markov chain Monte Carlo (MCMC) methods such as Metropolis adjusted Langevin algorithm or Hamiltonian Monte Carlo [36] and preconditioning (of gradients) techniques [17]. Such MCMC methods are known to scale with dimensionality in certain contexts (see, for example, [22, 36, 2, 26]) and may solve large-scale Bayesian problems (see, for example, [8, 28]). Sequential Monte Carlo (SMC) methods or other methods based on interacting clones/particles are an interesting alternative for Bayesian sampling. These methods have also shown to be stable in certain high-dimensional context [4], in particular adaptive SMC simulation [5], and have been employed to solve large-scale Bayesian inverse problems (see, for example, [7, 30]). However, in a general setting, it is difficult to argue theoretically for the superiority of one sampler over another. From an empirical point of view, studies tend to show that, while both methods achieve similar outcomes when the optimal solutions are not in the tail of the specified prior distribution, SMC’s strength lies in the dispersion of its interacting particles compared to a single point in MCMC sampling [29]. Let us mention that other sampling methods such as parallel tempering have been proposed to accelerate MCMC convergence [31]. Nevertheless, sampling with any of these methods becomes difficult in the present context, where only a too small budget of evaluations of the score function $S^{\ast}$ is affordable.
@@
In recent years, numerous MCMC and SMC methods based on reduced score $S$ approximating $S^{\ast}$ have been studied to reduce the simulation cost. Among MCMC approaches for inverse Bayesian problems, as expected, convergence results have been obtained for a sequence of reduced scores converging to $S^{\ast}$ [15]. Refinements of this approach involve integrating an iterative update of the $S^{\ast}-S$ approximation error into a reduced MCMC sampling scheme using $S$, given a predetermined estimate of the error (based on the evaluation of the error for a set of samples drawn from the prior) [9]. It should be noted, however, that a priori sampling to build the reduced model, characterize the approximation error or do both is probably not a relevant approach in our context. Indeed, in Bayesian inverse problems with informative measurements or in low-temperature problems, the target distribution usually concentrates its mass in a small and specific but unknown region of the prior support. More adapted to this context is the strategy to adapt the reduced model during the simulation, using samples whose distribution gradually matches the final target distribution.
@@
Adaptive construction of the reduced score embedded in a tempering scheme was suggested in [33]. As mentioned above, such an adaptive construction of the approximation is relevant to our context, but this work remains quite specific as it relies on parametric Gaussian proposals designed by a cross-entropy method [39]. Moreover, the adaptive construction of reduced score does not use possible error estimates.
@@
Regarding SMC approaches, multilevel SMC has been proposed to optimize the product cost $\times$ variance of the estimates; in the latter ‘multilevel’ refers to a sequence of approximations of the score function at different discretization levels [6]. However, multilevel SMC requires precise characterization of the model-dependent variance bounds of the estimates, which is a tedious task. In addition, although it relies on a given hierarchy of reduced models or approximations, these are not updated online using snapshots.
@@
An adaptive SMC approach using a predetermined hierarchy of reduced scores has also been proposed in [32]. This work tackles the tricky problem of selecting the right reduced model at the current temperature during simulation. Indeed, choosing a proper criterion to decide along the simulation when to refine the reduced score is quite an open question in general [24], although various heuristics have been proposed in the context or rare event simulation relying on some error quantification [3, 27, 13]. In their work, the authors of [32] propose a heuristic consisting of minimizing the variation of the particle weights over the two possible scenarios: refining or maintaining the reduced score. In this work, we go beyond these ideas and generally answer the question of when to update the reduced score and with which snapshots.
@@
The authors of [32] have also proposed an analogous method [40] in the context of rare event simulation, in which a predetermined hierarchy of reduced model is given. The authors of the present work have also proposed a method similar to the one studied in the present paper in a rare event context: the adaptive reduced multilevel [^1] Note that the term ‘multilevel’ in ARMS refers to the level sets of the score function [11], while in ‘multilevel SMC’ this term refers to the different precision levels of the simulation [25, 6] splitting (ARMS) algorithm [13]. ARMS is based on a sequential simulation scheme very similar to SMC, which adapts a reduced score approximation while getting closer to the target rare event distribution. This method is nevertheless specific to the rare event simulation context in the sense that it is dedicated to sample targets being the reference distribution conditionally to be in some rare event set. The method proposed in this paper is called adaptive reduced tempering (ART) and can be seen as a generalization of the ARMS algorithm to the case of general target distributions of the form of (1.1), which are relevant for Bayesian problems. However, as explained e.g. in Section 1.1, the ART method is also suited to the context of rare event simulation, by considering the latter as a zero temperature limit of a Gibbs family.
@@ 1.3 Contributions
Compared to the pieces of work mentioned in Section 1.2, the main novelty of the proposed ART algorithm is to propose an adaptive rule that triggers, within a SMC sampling method with tempered surrogate targets, when (using an entropy criterion) and where (using an importance distribution and an error quantification) to sample a new snapshot of the true model (serving to update the reduced model).
To do so, we consider the following exponential family of importance distributions
@@
where the normalization $Z_{\beta}\displaystyle\mathop{=}^{{\rm def}}\pi(\mathrm{e}^{\beta S})$ represents an approximation of $Z^{\ast}_{\beta}$. An important idea is to interpret the latter as proposal distributions (also called importance distributions), in the sense of importance sampling, of the target family (1.1) . In the context of rare event simulation, using proposal distributions in this exponential family rather than e.g. in the form of indicator of events $\propto\,{\rm l}\hskip-1.99168pt{\rm l}_{S\geq\ell}\mathrm{d}\pi$ with $\ell\in\mathbb{R}$ enabling to get rid of the possible nesting conditions between successive proposals in the algorithm (see [13]).
@@
where the reduced score $S^{(k)}$ is much cheaper to compute, and only the evaluation of the importance weight are expensive. Very importantly, as the iteration progresses, the proposal will adapt by adjusting the inverse temperature $\beta^{(k)}$ and refining the approximation of the reduced score $S^{(k)}$ using the history of the previous computations of the true model. In this framework, three issues emerge.
A first important one is that one has to address the definition of a criterion for setting the lowest inverse temperature $\beta^{(k)}$ suitable for the accuracy of the current reduced score approximation $S^{(k)}$. We recall the definition of the relative entropy or Kullback-Leibler divergence between probabilities defined by
$\mathrm{Ent}(\eta\mid\mu)\displaystyle\mathop{=}^{{\rm def}}\int\ln\frac{\mathrm{d}\eta}{\mathrm{d}\mu}\mathrm{d}\eta$
when $\eta$ is dominated by $\mu$ and extended to $\infty$ otherwise. We propose to constrain the relative entropy between a surrogate target denoted $\check{\eta}_{\beta}$ and the proposal as follows (for a given threshold $c_{1}$):
@@
Relative entropy can be used to estimate the log of the cost of importance sampling of the surrogate $\check{\eta}^{(k)}_{\beta^{(k)}}$ by the proposal $\mu^{(k)}_{\beta^{(k)}}$, as will be discussed in more detail in Section 2.2. This criteria ensures that the proposal can yield samples that are reasonably typical under the true target $\eta^{\ast}_{\beta^{(k)}}$. The surrogate target is constructed using the approximate error bound (1.2) of the reduced score approximation as follows:
@@
The choice is motivated by the fact that $\check{\eta}_{\beta}$ is roughly speaking a version of $\mu_{\beta}$ where typical states with large expected errors (hence ‘far away’ from the already sampled snapshots) are penalized.
A second issue that we will address in this work is the tuning of the expensive snapshot sampling and its associated reduced modeling update. At iteration $k+1$ the snapshot $X_{k+1}$ will be sampled from the proposal $\mu^{(k)}_{\beta^{(k)}}$ which ensures thanks to (1.5) a sampling in areas consistent with the true target distribution $\eta^{\ast}_{\beta^{(k)}}$ at inverse temperature $\beta^{(k)}$. In the first phase of the algorithm, we sample snapshots with an additional strong penalty weight favoring candidate states with a larger error (hence requiring priority correction). In the final phase of the algorithm, we remove that penalty to obtain snapshot-based optimal importance sampling estimation.
Refer to caption Refer to caption
Fig. 1: Illustration of one iteration of the main ART algorithm. In the first picture, a snapshot $X_{k+1}$ (cyan) has been sampled with an error-weighted version (black) of the proposal $\mu^{(k)}_{\beta^{(k)}}$ (red). In the second picture, this proposal (red) is updated with the new snapshot using the reduced model. In the third picture, temperature is decreased until the cross-entropy between the updated proposal $\mu^{(k+1)}_{\beta}$ (red) and the surrogate target $\check{\eta}^{(k+1)}_{\beta}$ (green) reaches the small threshold $c_{1}$. Iteration follows. Note that the surrogate target is concentrated on the typical states of the proposal that have relatively small error. The entropic criteria ensures matching between target and proposal.
The main two successive steps: i) temperature tuning, and ii) new snapshot sampling, do form the core steps of our methodology and are summarized in Fig. 1. A corresponding moderately detailed idealized pseudo-code can be found in Section 2.7 after more detailed descriptions of the various steps.
A more optional issue that has been addressed in this work is the simulation of a new proposal when at the end of iteration $k$ the reduced score is updated $S^{(k)}\to S^{(k+1)}$. We will propose to bridge proposals by searching among past simulations for an iteration index $k_{bridge}$ with associated approximate score $S^{(k_{bridge})}$, and by finding an inverse temperature $\beta_{bridge}$ such that the relative entropy between the new proposal $\mu_{\beta_{\mathrm{bridge}}}^{(k+1)}$ and the past proposal $\mu_{\beta^{(k_{\mathrm{bridge}})}}^{(k_{\mathrm{bridge}})}$} is kept below a certain standard threshold:
@@
The paper is organized as follows. An idealized version of the proposed algorithm, in which we assume the exact evaluation of expectations and the exact sampling of distributions, is presented in Section 2 together with the underlying methodological concepts. The proposed importance sampling estimator is shown to be consistent in such an idealized setting. We then provide in Section 3 a practical implementation of this idealized algorithm using SMC simulations. In Section 4, we exhibit by numerical experiments its practical convergence in the case of rare event simulation and inverse Bayesian problems. Most importantly, by evaluating the computational complexity in relation to the estimation error, we show that the use of reduced modeling (as compared to similar methods with only true model evaluation) significantly reduces the simulation cost while achieving a comparable accuracy. The final section deals with conclusions.